OpenStudy (anonymous):

A compact disc costs $11.95. How much do four of them cost if they are on sale for 10% off the orgiginal price? PWEASEEEEEEEEE HELP!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Its for a friend :)

OpenStudy (anonymous):

yeah, yeah-- sure it is ^_^

OpenStudy (anonymous):

okay i confess...its for me... BAHAHAHAHA kidding xD

OpenStudy (anonymous):

sale price = marked price - discount = 11.95 - 0.1*11.95 = 10.76 so four discs will cost 10.76 * 4 = 43.04

OpenStudy (anonymous):

This was the way i was saying it in chat: \[4(11.95)(.90) = 43.02\]

OpenStudy (anonymous):

The .90 is there because when you take 10% off 100%, there is 90% left.

OpenStudy (anonymous):

so if they are on a 10% discount, they are 90% of the original, right? \[11.95 * 0.9 = 10.755 \approx 10.76\]and there are 4 of em, so\[4*10.755 = 43.02\] But as Harkirat postet - the store would probably calculate it with the rounded 10.76 and thus you'd have to pay 2 cents more... :P